# Renormalization Group Evolution for In-medium Energy Correlators

### v1

@@ 1 Introduction
Among the broad class of jet substructure observables, the so-called energy-energy correlator (EEC) has emerged as a particularly clean probe Larkoski:2013eya. Historically it was first introduced in the context of $e^{+}e^{-}$ annihilation as a global event shape, where it provided some of the earliest QCD precision tests Basham:1978bw; Brown:1981jv. These early theoretical studies established the calculability of the EEC in perturbation theory. More recently, they were shown to be sensitive to transverse momentum dependent parton distribution and fragmentation functions in deep inelastic scattering at the Electron-Ion Collider (EIC) Li:2020bub; Li:2021txc. The EEC for jet substructure is defined as the angular correlation of the energy flow restricted inside the jet and is sensitive to both vacuum and medium-induced radiation patterns. Renewed interest in the EECs has led to detailed analyses of the observable in the collinear limit (kinematic regime suitable for jet substructure), where factorization theorems and resummation techniques beyond leading logarithmic accuracy have been established for both light Dixon:2019uzg; Lee:2022uwt and heavy flavor jets Craft:2022kdo; Holguin:2022epo, yielding quantitative insight into mass effects and the dynamics of QCD radiation. Moreover, these theoretical advances have directly enabled a broad program of EEC measurements across multiple experiments CMS:2024mlf; ATLAS:2023tgo; ALICE:2024dfl; STAR:2025jut, motivating further their role as powerful probes of QCD.
@@
These results are important for at least three reasons. First, in $e^{+}e^{-}$ and proton–proton collisions, the anomalous dimensions of the EEC have proven to be a robust feature, enabling the extraction of fundamental quantities such as the strong coupling constant. It is, therefore, natural to ask whether medium effects also induce corrections to the anomalous dimension, allowing the strength of jet–medium interactions to be extracted in a similarly robust manner. Second, the identification of a regime in which medium effects manifest solely as corrections to the anomalous dimension is of intrinsic interest, as it would indicate that the medium directly modifies the scale evolution of jets. This would provide a particularly transparent test of the picture of in-medium parton shower dynamics. Third, such a mechanism may also help explain the unexpectedly strong $p_{T}$ dependence observed in pp–Pb data Liang-Gilman:2025gjl, thereby offering a new perspective on experimental observations.
@@ 2.1 Energy Correlators for Semi-Inclusive Jets in the Vacuum
Figure 1: An illustration of the EEC of exclusive jets in the vacuum and after it is modified in a thin medium. In the asymptotic region, the EEC in the vacuum takes a power-law form with the time-like anomalous dimension $\gamma(\alpha_{s})=\sum_{n=0}\left(\frac{\alpha_{s}}{4\pi}\right)^{n+1}\gamma_{T}^{{\rm vac},(n)}$. It is also known that medium correction significantly modify the large-angle behavior of EEC, as shown by the red solid line. In this paper, we identify an important medium-induced resummation effect in the region $\Lambda_{\rm QCD}/p_{T}\ll\theta\ll 1/\sqrt{p_{T}L_{\rm eff}}$, resulting in a medium correction $\Delta\gamma(\alpha_{s})$ to the vacuum $\gamma(\alpha_{s})$. This is illustrated by the red dashed line.
@@ 2.1.3 The Full Next-to-Leading Order Result for $\Sigma_{q}$ in the Vacuum
$+\frac{\alpha_{s}(\mu^{2})}{2\pi}\left[\frac{\Theta(R-\theta)}{\theta^{2}}\frac{3C_{F}}{4}+\frac{295}{36}C_{F}\mathcal{J}_{\rm EEC,q}^{\rm vac,(0)}(\theta;p_{T},R,\mu)\right$. (14)
@@ 3.1 Scales in Nuclear Matter and Modified Energy Correlator Factorization Formula
Here, $m_{\rm eff}$ denotes the effective screening mass of the gluon propagator in the medium. In a high-temperature plasma, it can be estimated by the leading-order plasma Debye mass $m_{\rm eff}=m_{D}\sim g_{s}T$. Multiple scattering leads to transverse momentum broadening scale $\langle\delta k_{T}^{2}\rangle$, controlled by the product of the length of the medium $L$ and the jet transport parameter $\hat{q}\sim g_{s}^{4}T^{3}$ that characterizes the transverse momentum transfer squared per unit length. In a thin medium, the Landau–Pomeranchuk–Migdal (LPM) effect Landau:1953um; Migdal:1956tc introduces a new semi-hard scale $M_{\rm LPM}\sim\sqrt{2p_{T}/L}$ Ke:2023ixa; Ke:2024yzp. This can be understood as the typical virtuality of collinear fluctuations whose lifetime matches the medium length.
@@
Both of these cases can be achieved for the propagation of high-energy small $R$ jets in a thin (short length $L$) medium, such that $M_{\rm LPM}$ is the hardest among all the medium-related energy scales. For later convenience, we also define an angle related to the LPM effect [^2] Note that our estimates for $M_{\rm LPM}$ and $\theta_{\rm LM}$ are parametric, where we drop the explicit x(1x)x(1-x) kinematic dependence of the splitting which peaks at 1/4. Furthermore, the $M_{\rm LPM}$ scale is determined from the peak of an interference phase when the argument is $\sim\pi$. Thus, numerically more accurate values for those scales have an additional factor $\sqrt{4\pi}$. While irrelevant for our parametric discussion, this factor should be taken into account at least in $\theta_{\rm LPM}$ if one wishes to examine the in-medium EEC features in the phenomenological applications that follow.,
$\theta_{\rm LPM}=\sqrt{4\pi}{M_{\rm LPM}}/{p_{T}}=\sqrt{\frac{8\pi}{p_{T}L}}\,.$ (27)
@@ 3.4 TMD Splitting Functions at First Order in Opacity
Figure 2: Feynman diagrams that contributes to the TMD splitting functions at the amplitude level that are relevant for the calculation at opacity $N=1$. The red gluons are the Glauber gluon medium interactions.
@@
In a QGP medium, the effective screening mass $m_{\rm eff}$ is often taken to be the plasma Debye mass $m_{D}$, which can be perturbatively calculated at high temperature $m_{\rm eff}^{2}=m_{D}^{2}$. In cold nuclear matter $m_{\rm eff}\sim\Lambda_{\rm QCD}$. If we power expand in $m_{\rm eff}^{2}/Q^{2}$ with $Q$ being a harder scale in the problem, then the second term in $\int\frac{d^{2-2\epsilon}\mathbf{q}}{(2\pi)^{2-2\epsilon}}\frac{d\sigma_{FT}}{d^{2-2\epsilon}\mathbf{q}}\left[N_{q}(\mathbf{p}-\mathbf{q})-N_{q}(\mathbf{p})\right]$ becomes a scaleless integral, while the first term is regulated in DimReg.
@@ 4.1 Non-Contact Energy-Energy Correlator at First Order in Opacity
$\frac{dN^{RR}_{q\rightarrow q+g}}{dz^{+}dxd^{2-2\epsilon}\mathbf{k}d^{2-2\epsilon}\mathbf{q}}$ $=\frac{g_{s}^{2}C_{F}}{(2\pi)^{3-2\epsilon}}P_{qq,\epsilon}(x)\frac{1}{(2\pi)^{-2\epsilon}}\frac{\sum_{T}\rho_{T}^{-}(z^{+})g_{s}^{2}C_{T}/d_{A}}{(\mathbf{q}^{2}+m_{\rm eff}^{2})^{2}}g_{s}^{2}\left\{\frac{C_{F}}{[\mathbf{k}-(1-x)\mathbf{q}]^{2}}\right$.
@@
$\times g_{s}^{2}\left\{\frac{C_{A}}{(\mathbf{k}-x\mathbf{q})^{2}}+\frac{2C_{A}}{\mathbf{Q}_{3}^{2}}\Phi\left(\frac{(\mathbf{k}-\mathbf{q})^{2}z^{+}}{2x(1-x)P^{+}}\right)+\frac{2C_{A}}{\mathbf{Q}_{2}^{2}}\Phi\left(\frac{\mathbf{k}^{2}z^{+}}{2x(1-x)P^{+}}\right)\right$.
@@
$\times g_{s}^{2}\left\{\frac{C_{F}}{(\mathbf{k}-x\mathbf{q})^{2}}+\frac{2C_{F}}{\mathbf{Q}_{3}^{2}}\Phi\left(\frac{(\mathbf{k}-\mathbf{q})^{2}z^{+}}{2x(1-x)P^{+}}\right)+\frac{2C_{F}}{\mathbf{Q}_{2}^{2}}\Phi\left(\frac{\mathbf{k}^{2}z^{+}}{2x(1-x)P^{+}}\right)\right$.
@@
$g\rightarrow q\bar{q}$: real emission with real collisions
@@ 4.2 The Coulomb Logarithm in the Non-Contact Energy-Energy Correlator
where $Q_{\rm LPM}=\sqrt{x(1-x)}M_{\rm LPM}=\sqrt{2x(1-x)P^{+}/L^{+}}$. This way, the LPM phase is just like two propagators in the calculation and we have plenty of techniques to manipulate them. From the method of region analysis of the integral, we can also identify the complete Coulomb log as $\ln\frac{P^{+}/L^{+}}{m_{\rm eff}^{2}}$. This phase-space region of this Coulomb logarithm is shown in Figure 4. Thus, we can identify that for exponential medium, the non-contact medium-induced EEC is
@@
Figure 3: Left: numerical results for the non-contact contribution to the energy–energy correlator in the vacuum (black dashed line) and with medium correction at first order in opacity (black solid line) for $p_{T}=200~\mathrm{GeV}$. The “brick” medium has a constant temperature $T=0.4$ GeV with maximum length 22 fm and the medium-modified EEC shows the transition across the characteristic angular scales $\theta_{\mathrm{LPM}}$. Right: only plotting the medium-induced non-contact component ($R$ + RV channels) normalized by $\alpha_{s}^{2}\rho_{\mathrm{eff}}L_{\rm eff}^{3}\ln\frac{p_{T}}{L_{\rm eff}m_{\rm eff}^{2}}$, as motivated by the analytic calculation of exponentially decaying medium. Different lines correspond to the variations of the jet $p_{T}$, the effective path length, and the screening mass of the brick medium. They all collapses curves, corroborating the leading poser dependence on medium parameters and the presence of the Coulomb logarithm $\ln\frac{p_{T}}{L_{\rm eff}^{+}m_{\rm eff}^{2}}$.
@@
The appearance of the Coulomb logarithm at this order may suggests that when including higher-order corrections, the phase space corresponding to the soft-gluon emission induced by the Glauber gluon might be reggeized and lead to additional evolution for the calculation of the non-contact medium-induced EEC. We will leave this to a future study.
@@ 5.1 Renormalization and Evolution with Medium Correction
$\Delta_{ji}\gamma^{\rm med}(N)$ can be read off from the pole part of the $f_{ji}^{\rm med}$ function
@@
This means that for angles smaller than $\theta_{\rm LPM}=\sqrt{8\pi/p_{T}L_{\rm eff}}$ we first initialize the EEC at scale $\mu=\theta p_{T}$ with $\Sigma_{\rm vac}$, then perform the medium modified evolution from $\mu=\theta p_{T}$ to $\mu=M_{\rm LPM}$
@@ 5.2 The Medium-Modified Semi-Inclusive Jet Function
$\mathcal{J}_{\rm EEC,c}^{\rm med,(1)}(\theta;z_{J},p_{T},R,\mu)$ $=\sum_{(jk)}\int_{0}^{\infty}dz^{+}\int_{0}^{1}dx\int d^{2}\mathbf{k}\int d^{2}\mathbf{q}$
@@
The full expression can only be evaluated numerically, but using region analysis, we can obtain the most important leading contributions analytically. Because the scale $\sqrt{2p_{T}/L_{\rm eff}}$ is smaller than the jet scale $p_{T}R$, the medium-induced collinear radiations outside of the cone will be power suppressed, and the relevant modes that causes jet energy loss is the collinear soft radiations.
@@
$\mathcal{H}_{ab\rightarrow cX}\otimes\mathcal{J}^{(1)}=\int\frac{dz_{J}}{z_{J}}\mathcal{H}_{ab\rightarrow cX}\left(\frac{p_{T}}{z_{J}}\right)\mathcal{J}_{\rm EEC,c}^{\rm med,(1)}(\theta;z_{J},p_{T},R,\mu)\,$, (92)
first consider regions where $z_{J}=x_{j}=x$ or $z_{J}=x_{k}=1-x$ is not much smaller than unity. Then using the assumed hierarchy of energy scales $p_{T}R\gg\sqrt{2p_{T}/L}$, the radiative process outside of the cone must by power suppressed by additional power of $\frac{2p_{T}/L}{(p_{T}R)^{2}}$. Including the opacity small parameter, the final correction is only of order $\frac{\rho_{\rm eff}L}{2p_{T}/L}\times\frac{2p_{T}/L}{(p_{T}R)^{2}}=\frac{\rho_{\rm eff}L}{(p_{T}R)^{2}}$. This is expected to be an negligible effect as it is suppressed by two powers of $p_{T}$. Nevertheless, in the soft-gluon emission region where $x_{g}\rightarrow 0$, there is a region that contributes at the power $\frac{\rho_{\rm eff}L}{p_{T}/L}$. To see this, we can start from the jet cone algorithm and express $\mathbf{k}^{2}$ with the LPM scaled dimensionless integration variable
@@
$\mathcal{H}_{ab\rightarrow qX}\otimes\mathcal{J}_{q}^{(1)}\approx$ $\int\frac{dx}{x}\mathcal{H}_{ab\rightarrow qX}\left(\frac{p_{T}}{x}\right)\left[\frac{dN^{N=1}_{q\rightarrow q+g}}{dxd^{2}\mathbf{k}}\right]_{+}d^{2}\mathbf{k}\Theta\left(1-x<\tilde{\mathbf{k}}^{2}\frac{2p_{T}/L}{\left(2p_{T}\tan\frac{R}{2}\right)^{2}}\right)$
@@
Because the only scale other than the jet algorithm is the splitting-specific LPM scale $2x(1-x)p_{T}/L_{\rm eff}$, we approximate the $\mathbf{k}^{2}$ in the jet algorithm condition by $\mathbf{k}^{2}\sim 2(1-x)p_{T}/L_{\rm eff}$ under the soft gluon approximation. This way, the cone condition becomes a lower bound of $x$

### v2

@@ 1 Introduction
Among the broad class of jet substructure observables, the so-called energy-energy correlator (EEC) has emerged as a particularly clean probe Chen:2020vvp . Historically it was first introduced in the context of $e^{+}e^{-}$ annihilation as a global event shape, where it provided some of the earliest QCD precision tests Basham:1978bw ; Brown:1981jv . These early theoretical studies established the calculability of the EEC in perturbation theory. More recently, they were shown to be sensitive to transverse momentum dependent parton distribution and fragmentation functions in deep inelastic scattering at the Electron-Ion Collider (EIC) Li:2020bub ; Li:2021txc . The EEC for jet substructure is defined as the angular correlation of the energy flow restricted inside the jet and is sensitive to both vacuum and medium-induced radiation patterns. Renewed interest in the EECs has led to detailed analyses of the observable in the collinear limit (kinematic regime suitable for jet substructure), where factorization theorems and resummation techniques beyond leading logarithmic accuracy have been established for both light Dixon:2019uzg ; Lee:2022uwt and heavy flavor jets Craft:2022kdo ; Holguin:2022epo , yielding quantitative insight into mass effects and the dynamics of QCD radiation. Moreover, these theoretical advances have directly enabled a broad program of EEC measurements across multiple experiments CMS:2024mlf ; ATLAS:2023tgo ; ALICE:2024dfl ; STAR:2025jut , motivating further their role as powerful probes of QCD.
@@
These results are important for at least three reasons. First, in $e^{+}e^{-}$ and proton–proton collisions, the anomalous dimensions of the EEC have proven to be a robust feature, enabling the extraction of fundamental quantities such as the strong coupling constant. It is, therefore, natural to ask whether medium effects also induce corrections to the anomalous dimension, allowing the strength of jet–medium interactions to be extracted in a similarly robust manner. Second, the identification of a regime in which medium effects manifest solely as corrections to the anomalous dimension is of intrinsic interest, as it would indicate that the medium directly modifies the scale evolution of jets. This would provide a particularly transparent test of the picture of in-medium parton shower dynamics. Third, such a mechanism may also help explain the unexpectedly strong $p_{T}$ dependence observed in pp–Pb data Liang-Gilman:2025gjl , thereby offering a new perspective on experimental observations. In the future, the in-medium EEC approach developed here can be extended to heavy flavor Andronic:2015wma ; Boer:2024ylx ; Xing:2024yrb for example including recent insights into quarkonium production Copeland:2023wbu ; Copeland:2025osx ; Copeland:2025vop .
@@ 2.1 Energy Correlators for Semi-Inclusive Jets in the Vacuum
Figure 1: An illustration of the EEC of exclusive jets in the vacuum and after it is modified in a thin medium. In the asymptotic region, the EEC in the vacuum takes a power-law form with the time-like anomalous dimension $\gamma(\alpha_{s})=\sum_{n=0}\left(\frac{\alpha_{s}}{4\pi}\right)^{n+1}\gamma_{T}^{{\rm vac},(n)}$. It is also known that medium correction significantly modify the large-angle behavior of EEC, as shown by the red solid line. In this paper, we identify an important medium-induced resummation effect in the region $\Lambda_{\rm QCD}/p_{T}\ll\theta\ll\sqrt{8\pi/p_{T}L_{\rm eff}}$, resulting in a medium correction $\Delta\gamma(\alpha_{s})$ to the vacuum $\gamma(\alpha_{s})$. This is illustrated by the red dashed line.
@@ 2.1.3 The Full Next-to-Leading Order Result for $\Sigma_{q}$ in the Vacuum
$+\frac{\alpha_{s}(\mu^{2})}{2\pi}\left[\frac{\Theta(R-\theta)}{\theta^{2}}\frac{3C_{F}}{4}+\frac{295}{36}C_{F}\mathcal{J}_{\rm EEC,q}^{\rm vac,(0)}(\theta;p_{T},R,\mu)\right$.
@@ 3.1 Scales in Nuclear Matter and Modified Energy Correlator Factorization Formula
Here, $m_{\rm eff}$ denotes the effective screening mass of the gluon propagator in the medium. In a high-temperature plasma, it can be estimated by the leading-order plasma Debye mass $m_{\rm eff}=m_{D}\sim g_{s}T$. Multiple scattering leads to transverse momentum broadening scale $\langle\delta k_{T}^{2}\rangle$, controlled by the product of the length of the medium $L$ and the jet transport parameter $\hat{q}\sim g_{s}^{4}T^{3}$ that characterizes the transverse momentum transfer squared per unit length. In a thin medium, the Landau–Pomeranchuk–Migdal (LPM) effect Landau:1953um ; Migdal:1956tc introduces a new semi-hard scale $M_{\rm LPM}\propto\sqrt{p_{T}/L}$ Ke:2023ixa ; Ke:2024yzp . This can be understood as the parametric scale of the typical virtuality of collinear fluctuations whose lifetime matches the medium length.
@@
Both of these cases can be achieved for the propagation of high-energy small $R$ jets in a thin (short length $L$) medium, such that $M_{\rm LPM}$ is the hardest among all the medium-related energy scales. For later convenience, we also define an angle related to the LPM effect [^2] Note that our estimates for $M_{\rm LPM}$ and $\theta_{\rm LPM}$ are parametric, where we drop the explicit x(1x)x(1-x) kinematic dependence. For $M_{\rm LPM}$, this adequately describes its power and logarithmic appearance. If one wishes to examine the in-medium EEC features in the phenomenological applications that follow, e.g. transitional behavior around $\theta_{\rm LPM}$, a more accurate estimate of $\theta_{\rm LPM}$ is desirable. From the first peak of the LPM interference factor $1-\cos\frac{\mathbf{k}^{2}L}{2x(1-x)p_{T}}$, the argument should equal $\pi$. Taking the peak value of x(1x)=1/4x(1-x)=1/4 in $\theta=|{\bf k}|/[x(1-x)p_{T}]$, its is easy to obtain Eq. (26). ,
$\theta_{\rm LPM}=\sqrt{\frac{8\pi}{p_{T}L}}\,.$ (26)
@@ 3.4 TMD Splitting Functions at First Order in Opacity
Figure 2: Feynman diagrams that contribute to the TMD splitting functions at the amplitude level that are relevant for the calculation at opacity $N=1$. The red gluons are the Glauber gluon medium interactions.
@@
Here, $C_{T}$ is the quadratic Casimir of the target parton representation and $d_{A}=8$ is the dimension of the adjoint representation. In a QGP medium, the effective screening mass $m_{\rm eff}$ is often taken to be the plasma Debye mass $m_{D}$, which can be perturbatively calculated at high temperature $m_{\rm eff}^{2}=m_{D}^{2}$. In cold nuclear matter $m_{\rm eff}\sim\Lambda_{\rm QCD}$. If we power expand in $m_{\rm eff}^{2}/Q^{2}$ with $Q$ being a harder scale in the problem, then the second term in $\int\frac{d^{2-2\epsilon}\mathbf{q}}{(2\pi)^{2-2\epsilon}}\frac{d\sigma_{FT}}{d^{2-2\epsilon}\mathbf{q}}\left[N_{q}(\mathbf{p}-\mathbf{q})-N_{q}(\mathbf{p})\right]$ becomes a scaleless integral, while the first term is regulated in DimReg.
@@ 4.1 Non-Contact Energy-Energy Correlator at First Order in Opacity
$\frac{dN^{RR}_{q\rightarrow q+g}}{dz^{+}dxd^{2-2\epsilon}\mathbf{k}d^{2-2\epsilon}\mathbf{q}}=$ $\frac{g_{s}^{2}C_{F}}{(2\pi)^{3-2\epsilon}}P_{qq,\epsilon}(x)\frac{1}{(2\pi)^{2-2\epsilon}}\frac{\sum_{T}\rho_{T}^{-}(z^{+})g_{s}^{2}C_{T}/d_{A}}{(\mathbf{q}^{2}+m_{\rm eff}^{2})^{2}}$
$\times g_{s}^{2}\left\{\frac{C_{F}}{[\mathbf{k}-(1-x)\mathbf{q}]^{2}}\right$.
@@
$\times g_{s}^{2}\left\{\frac{C_{A}}{(\mathbf{k}-x\mathbf{q})^{2}}\right$.
$+\frac{2C_{A}}{(\mathbf{k}-\mathbf{q})^{2}}\Phi\left(\frac{(\mathbf{k}-\mathbf{q})^{2}z^{+}}{2x(1-x)P^{+}}\right)+\frac{2C_{A}}{\mathbf{k}^{2}}\Phi\left(\frac{\mathbf{k}^{2}z^{+}}{2x(1-x)P^{+}}\right)$
@@
$\times g_{s}^{2}\left\{\frac{C_{F}}{(\mathbf{k}-x\mathbf{q})^{2}}\right$.
$+\frac{2C_{F}}{(\mathbf{k}-\mathbf{q})^{2}}\Phi\left(\frac{(\mathbf{k}-\mathbf{q})^{2}z^{+}}{2x(1-x)P^{+}}\right)+\frac{2C_{F}}{\mathbf{k}^{2}}\Phi\left(\frac{\mathbf{k}^{2}z^{+}}{2x(1-x)P^{+}}\right)$
@@
$g\rightarrow q\bar{q}$: real emission with virtual collisions
@@ 4.2 The Coulomb Logarithm in the Non-Contact Energy-Energy Correlator
where $Q_{\rm LPM}=\sqrt{2x(1-x)P^{+}/L^{+}}$. This way, the LPM phase is just like two propagators in the calculation and we have plenty of techniques to manipulate them. From the method of region analysis of the integral, we can also identify the complete Coulomb log as $\ln\frac{P^{+}/L^{+}}{m_{\rm eff}^{2}}$. This phase-space region of this Coulomb logarithm is shown in Figure 4. Thus, we can identify that for exponential medium, the non-contact medium-induced EEC is
@@
Figure 3: Left: numerical results for the non-contact contribution to the energy–energy correlator in the vacuum (black dashed line) and with medium correction at first order in opacity (black solid line) for $p_{T}=200~\mathrm{GeV}$. The “brick” medium has a constant temperature $T=0.4$ GeV with maximum length 22 fm and the medium-modified EEC shows the transition across the characteristic angular scales $\theta_{\mathrm{LPM}}$, which is defined in Eq. (26). Right: only plotting the medium-induced non-contact component ($R$ + RV channels) normalized by $\alpha_{s}^{2}\rho_{\mathrm{eff}}L_{\rm eff}^{3}\ln\frac{p_{T}}{L_{\rm eff}m_{\rm eff}^{2}}$, as motivated by the analytic calculation of exponentially decaying medium. Different lines correspond to the variations of the jet $p_{T}$, the effective path length, and the screening mass of the brick medium. They all collapse to a universal curve, corroborating the leading power dependence on medium parameters and the presence of the Coulomb logarithm $\ln\frac{p_{T}}{L_{\rm eff}^{+}m_{\rm eff}^{2}}$.
@@
The appearance of a Coulomb logarithm at this order suggests that, once higher-order corrections are included, the phase space associated with soft-gluon emission induced by Glauber exchange may become Reggeized, leading to additional evolution effects in the calculation of the non-contact medium-induced EEC. We leave a detailed investigation of this possibility to future work.
@@ 5.1 Renormalization and Evolution with Medium Correction
$\Delta\gamma_{ji}^{\rm med}(N)$ can be read off from the pole part of the $f_{ji}^{\rm med}$ function
@@
This means that for angles parametrically smaller than $\theta_{\rm LPM}$ we first initialize the EEC at scale $\mu=\theta p_{T}$ with $\Sigma_{\rm vac}$, then perform the medium modified evolution from $\mu=\theta p_{T}$ to $\mu=M_{\rm LPM}$
@@ 5.2 The Medium-Modified Semi-Inclusive Jet Function
$\mathcal{J}_{\rm EEC,c}^{\rm med,(1)>}(\theta;z_{J},p_{T},R,\mu)=$ $\sum_{(jk)}\int_{0}^{\infty}dz^{+}\int_{0}^{1}dx\int d^{2}\mathbf{k}\int d^{2}\mathbf{q}$
@@
The full expression can only be evaluated numerically, but using region analysis, we can obtain the most important leading contributions analytically. Because the scale $M_{\rm LPM}$ is smaller than the jet scale $p_{T}R$, the medium-induced collinear radiations outside of the cone will be power suppressed, and the relevant modes that causes jet energy loss is the collinear soft radiations.
@@
$\mathcal{H}_{ab\rightarrow cX}\otimes\mathcal{J}^{(1)>}=\int\frac{dz_{J}}{z_{J}}\mathcal{H}_{ab\rightarrow cX}\left(\frac{p_{T}}{z_{J}}\right)\mathcal{J}_{\rm EEC,c}^{\rm med,(1)>}(\theta;z_{J},p_{T},R,\mu)\,$, (91)
first consider regions where $z_{J}=x_{j}=x$ or $z_{J}=x_{k}=1-x$ is not much smaller than unity. Then using the assumed hierarchy of energy scales $p_{T}R\gg M_{\rm LPM}=\sqrt{2p_{T}/L}$, the radiative process outside of the cone must by power suppressed by additional power of $\frac{p_{T}/L}{(p_{T}R)^{2}}$. Including the opacity small parameter, the final correction is only of order $\frac{\rho_{\rm eff}L}{p_{T}/L}\times\frac{p_{T}/L}{(p_{T}R)^{2}}=\frac{\rho_{\rm eff}L}{(p_{T}R)^{2}}$. This is expected to be an negligible effect as it is suppressed by two powers of $p_{T}$. Nevertheless, in the soft-gluon emission region where $x_{g}\rightarrow 0$, there is a region that contributes at the power $\frac{\rho_{\rm eff}L}{p_{T}/L}$. To see this, we can start from the jet cone algorithm and express $\mathbf{k}^{2}$ with the LPM scaled dimensionless integration variable
@@
$\mathcal{H}_{ab\rightarrow qX}\otimes\mathcal{J}_{q}^{(1)>}\approx$ $\int\frac{dx}{x}\mathcal{H}_{ab\rightarrow qX}\left(\frac{p_{T}}{x}\right)\int d^{2}\mathbf{k}\frac{dN^{N=1}_{q\rightarrow q+g}}{dxd^{2}\mathbf{k}}\Theta\left(1-x<\tilde{\mathbf{k}}^{2}\frac{2p_{T}/L}{\left(2p_{T}\tan\frac{R}{2}\right)^{2}}\right)$
@@
Because the only scale other than the jet algorithm is the splitting-specific LPM scale squared $2x(1-x)p_{T}/L_{\rm eff}$, we approximate the $\mathbf{k}^{2}$ in the jet algorithm condition by $\mathbf{k}^{2}\sim 2(1-x)p_{T}/L_{\rm eff}$ under the soft gluon approximation. This way, the cone condition becomes a lower bound of $x$